Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 12,750 $ 16,650 $ 29,400
Estimated variable manufacturing overhead per machine-hour $ 2.50 $ 3.30
Job P Job Q
Direct materials $ 24,000 $ 13,500
Direct labor cost $ 29,800 $ 11,900
Actual machine-hours used:
Molding 2,800 1,900
Fabrication 1,700 2,000
Total 4,500 3,900

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

1. What was the company’s plantwide predetermined overhead rate?

2. How much manufacturing overhead was applied to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

3. What was the total manufacturing cost assigned to Job P? (Do not round intermediate calculations.)

4. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

5. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.)
Answer #1

The requirement is to compute plantwide overhead rate. That means both the molding and fabrication department data needs to be combined to compute plantwide overhead rate.

1.

Variable manufacturing overhead = Rate X machine hours

Molding = 2.5 * 2500 = $ 6,250

Fabrication = 3.3 * 1500 = $ 4,950

Variable manufacturing overhead = 6250 + 4950 = $ 11,200

Total manufacturing overhead = 29,400 + 11,200 = $ 40,600

Hence, plantwide overhead rate = Total estimated manufacturing overhead/estimated machine hours

                                                                = 40,600 / 4,000 = $ 10.15 per hour

2.

Job P

Job Q

Machine hours (actual)

4,500

3,900

X    Overhead rate

10.15

10.15

Overheads applied

$ 45,675

$ 39,585

3.

Job P

Direct materials

24,000

Direct labor cost

29,800

Overheads applied

45,675

total manufacturing cost

$ 99,475

4.

Job P

total manufacturing cost

$ 99,475

Number of Units

20

Unit product cost

$ 4973.75 per unit

5.

Job Q

Direct materials

13,500

Direct labor cost

11,900

Overheads applied

39,585

total manufacturing cost

$ 64,985